# Visual Writing: Writing by Manipulating Visual Representations of Stories

### v1

@@ Abstract.
We introduce “visual writing”, an approach to writing stories by manipulating visuals instead of words. Visual writing relies on editable visual representations of time, entities, events, and locations to offer representations more suited to specific editing tasks. We propose a taxonomy for these representations and implement a prototype software supporting the visual writing workflow. The system allows writers to edit the story by alternating between modifying the text and manipulating visual representations to edit entities, actions, locations, and order of events. We evaluate this workflow with eight creative writers and find visual writing can help find specific passages, keep track of story elements, specify edits, and explore story variations in a way that encourages creativity.
@@ 1. Introduction
Writing stories and exploring alternative narratives can be challenging. Although using large language models (LLMs) can help writers (Chakrabarty et al., 2024), it remains difficult for users to formulate their intent using words (Dang et al., 2022b) and review the generated text to ensure it aligns with their intent (Masson et al., 2024).
@@
Consider a task as simple as changing the location of a character in a story: the cat no longer goes to the barn but instead wanders about the lake. Using “replace all” instances of ‘barn’ is insufficient because substituting a location comes with implications. For example, descriptions and actions pertaining to the barn would need to be updated to reflect the new context of being by the lake. Further, not all instances necessarily need to be impacted. The fact that the cat no longer visits the barn should not result in the barn being completely removed from the story. Instead, writers must carefully plan to find all and only the references to the location associated with the character’s change of action and make tedious revisions to the text to maintain spatial and semantic consistencies. This busywork is prone to errors, and writers often resort to creating external documents such as maps, spreadsheets, and timelines that they painstakingly update to keep track of the characters, events, and locations (Marshall, 2014; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014). While LLMs could help with this busywork, expressing distinct changes can be challenging with prompts (Masson et al., 2024), and the extent and suitability of the changes made by the generative model could be difficult to verify.
@@
The challenge stems from the mismatch between the reasoning and representation spaces: in our above example, the writer performs spatial reasoning using a textual representation. A spatial representation, such as a map of the story world, would be more suited for such a task, where changing the location of the character becomes as simple as moving its position on the map. In text, edits relevant to space and locations in the story are comparatively challenging to accomplish, which can be a deterrent for writers wanting to change the narrative and explore variations. And this same argument applies to all other aspects of a story such as time, events, and characters, for which we posit superior representations can be designed.
@@
We define visual writing as the use of visual representations of story elements as an input medium for expressing writing intents. For each story element, we explore the use of a visual representation that conceptually aligns with that element. Depending on their editing task, users can decide to edit the text of the story, in which case the visual representations are updated to match the new text. Alternatively, they can manipulate the visual representations through direct manipulation to suggest changes to the text and rapidly explore variations. We discuss a taxonomy of these visual representations, drawing on the story elements framework from narratology (§3). In a prototype software (§4), we implement three of these visual representations: (1) a diagrammatic view of the entities in the story and their relationships, allowing to add, remove, and edit characters and add new actions between them (Figure 1.c); (2) a spatial view of the locations in the story and the locations of the entities, allowing to add new locations and move entities (Figure 1.d); and (3) a timeline of the events in the story, allowing to reorder the events, quickly find specific scenes, and precisely modify the entities and locations of the selected scenes by leveraging the other two views (Figure 1.b).
We validate this approach in a user study with eight creative writers (§5). Results show participants appreciated the visual representations to locate scenes and keep track of story elements while writing. They were successful in expressing their editing intents through direct manipulations of the visual elements. Furthermore, participants appreciated this approach to edit text from an exploration and creativity perspective. This points to the potential of visual representations for writing and calls for more work in that direction.
@@ 2. Related Work
Visual writing draws inspiration from work on visualizing stories and efforts to give alternative representations to edit content. Below, we review the literature on these topics.
@@ 2.1. Visualizing Stories
Visual representations of stories and story patterns have a long-standing presence in writing, with staple examples including Freytag’s pyramid story (Freytag, 1895) and abstract diagrams capturing the “shape of the story” (Pope, 1998; Comberg, [n. d.]; Reagan et al., 2016). Such visual representations are not only useful for teaching the craft of storytelling, they are also considered as useful aids to support the writing process. Referring to the visual representations, screenplay writer E. Williams says: “I find story shape to be a great tool […] to make sure my story holds together” (Williams, [n. d.]).
@@
There exist many techniques to visualize stories, spanning from classical charts like word clouds, to more artistic visualizations like calligrams (Brath, 2021). Often, these visualizations summarize entire stories in an attempt to highlight a specific facet (e.g., emotions (Maharjan et al., 2018) or topic recurrence (Zhu, 2013)). For example, Storyline visualizations (Munroe, 2009; Tanahashi and Kwan-Liu Ma, 2012) depict the evolution of the relationships between characters through timelines. StoryCake (Qiang and Bingjie, 2016) uses a hierarchical plot to visualize the structure of nonlinear stories. Story Curves (Kim et al., 2018) visualizes non-linear narratives by showing both the order in which the events are told and the order of the underlying story. StoryPrint (Watson et al., 2019) shows the scenes and character emotions through circular timelines. Each of these visual representations presents the ability to support readers, writers, and analysts alike in appreciating or even analyzing the stories from a specific lens.
@@
Other story visualizations have been used in tools to help during script and story writing. For example, CARDINAL (Marti et al., 2018) is a tool to help screenwriters visualize their script through different views, including a visualization of the interactions between characters and an animated 2D and 3D views to play the scene and visualize locations and movements of the characters. Similarly, Portrayal (Hoque et al., 2023) attempts to help writers develop their characters by extracting and presenting information related to the characters in interactive visualizations.
Our work explores the potential of expanding the role of such visualizations to also be manipulable. That is, by mapping visualization elements directly to entities, events, and locations in the text story, we explore using visualizations as the medium to modify the story. This approach is motivated by the potential of visual representations to express intent, which we discuss next.
2.2. Editing Content Through Alternative Representations
@@
Programming might be the domain that most embraced the use of alternative representations. Approaches include visual programming (Myers, 1986; Burnett and McIntyre, 1995), which allows generating programs by manipulating elements in more than one dimension, code projections (Gobert and Beaudouin-Lafon, 2023; Ko and Myers, 2006) which augment programming languages with tools and widgets that are alternative representations of fragments of code, and bi-directional programming (Mayer et al., 2018; Hempel et al., 2019) where the output of a program can be manipulated to edit the program itself. Similar to our work, the motivation behind these approaches is that some tasks are easier done using different representations: it is easier to find a colour using a colour picker than by typing its hexadecimal code, and it is easier to move an object on the screen through direct manipulation (Shneiderman, 1983) than by typing its coordinates.
@@
In the creative writing domain, this idea of manipulating projections or visual representations to edit stories is largely untapped, which is surprising given the reliance on visual representations to capture story concepts (Freytag, 1895; Comberg, [n. d.]). Instead, most writing support tools ask users to manipulate text using text, typically through prompts or suggestions (Calderwood et al., 2020; Buschek et al., 2021; Yuan et al., 2022; Dang et al., 2023; Lee et al., 2024; Reza et al., 2024). There are some exceptions, such as the proposal of Kempen et al. (1987) to represent and manipulate text in different ways, such as manipulating the syntax tree to allow restructuring sentences. Similarly, Arnold et al. (2021) suggested that words could be directly manipulated to be reordered. Beyond syntactic level modifications, Dang et al. (2022a) proposed a system that allows reorganizing, merging, and removing paragraphs by manipulating textual summaries instead of the text. In the same vein, VISAR (Zhang et al., 2023) automatically generates an interactive visual outline of an argumentative text, which can be edited to change the structure of the argument. Finally, TaleBrush (Chung et al., 2022) leverages a canvas where sketching allows modulating the fortune of characters in a story.
@@
Our work contributes to this line of research by formalizing the concept of visual writing. The concept is rooted in the same principles as visual programming and code projections, positing that elements of a story become easier to understand and manipulate when presented in alternative forms. Specifically, we design visualizations grounded in narratology to explore this approach.
3. Taxonomy of Visual Representations
Narratology separates the chronological sequence of the story (fabula) from the narrative sequence (syuzhet) of the story. At its core, a fabula consists of four main story elements: events, actors111In narratology, actors refer to “agents that perform actions. They are not necessarily human” (Bal, 1997). To avoid confusion, we use the term entity instead throughout this paper., time, and location (Bal, 1997). Similarly, we define visual writing as the use of visuals that represent and manipulate these four elements.
In their simplest form, visual representations can depict elements one at a time. For example, a list of locations would allow adding, removing, and editing the characteristics of each location. And a list of entities, similar to the character sheets writers often create to plan their stories (SaveTheCat, 2014), would allow adding and removing entities. However, many important aspects of stories are encoded through the interaction between elements (e.g., the time of an event or the location of an entity). Below, using the narratology lens, we discuss a non-exhaustive taxonomy of possible visual representations as an initial guide to inform the design of tools supporting visual writing.
Spatial representations depict elements based on their location. For example, a spatial representation could be a map with dots, where each dot represents where events happened and where entities are located. Or it could be more abstract and only represent distances between elements. Spatial representations are used by systems like CARDINAL (Marti et al., 2018) to show the relative positioning of characters within a scene.
Event-driven representations depict elements based on the events they relate to. For example, a node graph could be an event-driven representation if each node represents an event and the edges encode another element such as the entities (in which case, connected nodes would represent events affecting the same entity) or time (in which case the directed edges would indicate the flow of the events). Some examples of event-driven representations are lists of outline points and event graphs.
Entity-driven representations depict elements based on the entities they relate to. For example, a node graph could be an entity-driven representation if each node represents an entity and the edges encode another element such as the events (in which case, each edge is an event connecting the entities involved in the event) or time (in which case, the directed edges indicate the order in which the entities intervene in the story). Entity-driven representations, such as character sheets and family trees, are often created by writers to keep track of character details (SaveTheCat, 2014).
Temporal representations depict elements based on the time they happen in the fictional world of the story or in the order they are presented in the text. For example, a timeline could arrange elements such as events (in which case, the events would be positioned based on when they occur) or entities (in which case, entities would be positioned based on when they appear or when they perform an action). Temporal representations have been more commonly used by previous work on visualizing stories such as Storylines (Munroe, 2009), StoryCake (Qiang and Bingjie, 2016), and StoryPrint (Watson et al., 2019).
The design space for the above categories is vast, and there exist many visualizations to bear inspiration from. Our goal is not to argue which specific design is best—we leave it to future work to explore different variations within these categories (and beyond). Instead, our work focuses on the potential of visual writing, i.e. the use of alternative representation of stories, as an input medium for expressing intent.
4. The Visual Writing Prototype
@@
To test the potential of visual writing, we developed a prototype system, shown in Figure 1, implementing three visual representations: a diagrammatic view of the entities and actions (Entity–Event), a spatial view of the locations and entities (Location–Entity), and a timeline of the events (Time–Event). The focus of the prototype was to help writers explore variations and perform spatial and temporal edits via visual writing. As such, the system allowed specifying edits through direct manipulations of the graphical elements (Shneiderman, 1982; Hutchins et al., 1985) and followed the principles to design creativity support tools (Shneiderman et al., 2006). Below, we detail the design and features of the system. All the examples and screenshots were taken using the system on the first three paragraphs of the story Alice’s Adventures in Wonderland by Lewis Carroll.

### v2

# Visual Story-Writing: Writing by Manipulating Visual Representations of Stories

@@ Abstract.
We define “visual story-writing” as using visual representations of story elements to support writing and revising narrative texts. To demonstrate this approach, we developed a text editor that automatically visualizes a graph of entity interactions, movement between locations, and a timeline of story events. Interacting with these visualizations results in suggested text edits: for example, connecting two characters in the graph creates an interaction between them, moving an entity updates their described location, and rearranging events on the timeline reorganizes the narrative sequence. Through two user studies on narrative text editing and writing, we found that visuals supported participants in planning high-level revisions, tracking story elements, and exploring story variations in ways that encourage creativity. Broadly, our work lays the foundation for writing support, not just through words, but also visuals.
@@ 1. Introduction
Creative writers manage many moving parts, from character arcs and causal chains to spatial coherence and narrative timing (Bal, 1997; Norton, 2011). Keeping track of everything is challenging, and it gets even more complicated when experimenting with story ideas (Zhao et al., 2025). Modifications require careful planning and multiple rounds of edits.
@@
Consider a task as simple as changing the location of a character in a story: the cat no longer goes to the barn but instead wanders about the lake. Simply replacing all instances of “barn” would cause inconsistencies unless descriptions and actions pertaining to the barn are also updated to reflect the new context of the lake. Moreover, the fact that the cat wanders about the lake does not mean that the barn needs to be removed entirely from the story. Instead, before revising the text, writers must identify references to the barn and tediously review the other characters’ locations to maintain spatial and semantic consistency. This busywork is prone to errors, and writers often resort to creating external documents, such as maps, spreadsheets, and timelines, that they painstakingly update to keep track of characters, events, and locations (Zhao et al., 2025; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014; Articy, 2014). Large language models (LLMs) could help, but expressing specific intents with a prompt is difficult (Masson et al., 2024; Zamfirescu-Pereira et al., 2023), and trying to explain which barn and which cat to edit will leave much room for misinterpretation. Further, the suitability of the changes made by the generative model could be difficult to verify.
@@
This challenge stems from the mismatch between the reasoning and representation spaces: in our barn example, the writer does spatial reasoning with a textual representation. A better approach would be to use a spatial representation, such as a map of the story world, where changing the location of a character becomes as simple as moving it on the map. This same argument applies to all other aspects of a story, such as time, events, and characters, for which we posit superior representations can be designed.
Therefore, we propose using representations that match the reasoning space. We define visual story-writing as the use of visual representations of story elements as both a reviewing tool and an input medium for expressing writing intents (Figure 1). This complements the writing workflow of writers with various experience levels. For example, an experienced writer may prefer editing the story’s text, in which case the visual representations updates, offering a visualization to help the writer track story elements. Alternatively, a beginner writer could manipulate the visual representations through direct manipulation to explore story variations.
@@
We articulate a framework grounded in narratology to design new visual representations of stories (§3). Applying our framework, we explore the potential of visual story-writing through a prototype software (§4) with three illustrative visual representations: (1) a diagrammatic view of the entities in the story and their relationships, allowing to add, remove, and edit characters and add new actions between them; (2) a spatial view of the entities in the story, allowing to add new locations and move entities; and (3) a timeline of the events in the story, allowing to reorder the narrative, quickly find specific scenes, and precisely modify the entities and locations of the selected scenes by leveraging the other views.
We validate this approach through two user studies looking at the different components of the writing process (Flower and Hayes, 1981). Results from our first study revealed that visuals helped plan high-level revisions, search, and promote critical reflection over having only text. Results from our second study show the potential of visual story-writing to express editing intents and help writers explore story variations in a way that supports creative expression. This points to the potential of visual story-writing and encourages further exploration.
@@ 2. Related Work
Visual story-writing draws inspiration from work on visualizing stories and efforts to give alternative representations to edit content.
@@ 2.1. Visualizing Stories
Visual representations of stories and story patterns have a long-standing presence in writing, with staple examples including Freytag’s pyramid story (Freytag, 2004) and abstract diagrams capturing the “shape of the story” (Pope, 1998; Comberg, [n. d.]; Reagan et al., 2016). Such visual representations are not only useful for teaching the craft of storytelling but are also considered useful aids to support the writing process. Referring to the visual representations, screenplay writer E. Williams says: “I find story shape to be a great tool […] to make sure my story holds together” (Williams, 2018).
@@
There exist many techniques to visualize stories, spanning from classical charts, like word clouds, to more artistic visualizations like calligrams (Brath, 2021). Often, these visualizations attempt to highlight a specific facet of the story (e.g., emotions (Maharjan et al., 2018), motion (Chung et al., 2025), fortune (Comberg, [n. d.]; Chung et al., 2022), or topic recurrence (Zhu, 2013)). For example, Storyline visualizations (Munroe, 2009; Tanahashi and Kwan-Liu Ma, 2012) depict the evolution of the relationships between characters through timelines. StoryCake (Qiang and Bingjie, 2016) uses a hierarchical plot to visualize the structure of nonlinear stories. Story Curves (Kim et al., 2018) visualizes non-linear narratives by showing both the order in which the events are told and the story’s chronological order. StoryPrint (Watson et al., 2019) shows the scenes and character emotions through circular timelines. Each of these visual representations presents the ability to support readers, writers, and analysts alike in appreciating or even analyzing the stories from a specific lens.
@@
Other story visualizations have been used in tools to help with script and story writing. For example, CARDINAL (Marti et al., 2018) helps screenwriters visualize their script through different views, including a visualization of the interactions between characters and animated 2D and 3D views to playout the scene and replay character movements. Similarly, Portrayal (Hoque et al., 2023) attempts to help writers develop their characters by extracting and presenting information related to the characters in interactive visualizations. Our work seeks to inform the design of story visualizations by proposing a new framework. This framework helps generate representations that might serve as a medium to edit the story, a topic we discuss next.
2.2. Editing Using Alternative Representations
Visualizations are increasingly used to not only encode data but also collect and modify new data (Bressa et al., 2024). In our work, visual representations can be edited to modify the narrative text in addition to exploring and analyzing it, making it an input visualization.
@@ 2.2. Editing Using Alternative Representations
One domain which has embraced the use of input visualizations is computer programming. For example, visual programming (Myers, 1986; Burnett and McIntyre, 1995) allows generating programs by manipulating elements in more than one dimension, code projections (Gobert and Beaudouin-Lafon, 2023; Ko and Myers, 2006) augments programming languages with tools and widgets that are alternative representations of fragments of code, and bi-directional programming (Mayer et al., 2018; Hempel et al., 2019) allows manipulating the output of a program to edit the program itself. Similar to our work, the motivation behind these approaches is that some tasks are easier done using different representations: it is easier to find a colour using a colour picker than by typing its hexadecimal code, and it is easier to move an object through direct manipulation (Shneiderman, 1983) than by typing coordinates.
@@
In the creative writing domain, the idea of manipulating projections or visual representations to edit stories is largely untapped despite the reliance on visual representations to capture story concepts (Freytag, 2004; Comberg, [n. d.]; Zhao et al., 2025). Instead, most writing support tools ask users to manipulate narratives using text, typically through prompts or suggestions (Calderwood et al., 2020; Buschek et al., 2021; Yuan et al., 2022; Dang et al., 2023; Lee et al., 2024; Reza et al., 2024). Some exceptions include the proposal of Kempen et al. (1987) to manipulate the syntax tree of a sentence to restructure it or the proposal to directly drag the words to reorder them (Arnold et al., 2021). Beyond syntactic level modifications, Dang et al. (2022) proposed a system that allows reorganizing, merging, and removing paragraphs by manipulating textual summaries. Moving even further from text as an input medium, the machine learning task of “visual storytelling” consists in generating a story from a sequence of images (Huang et al., 2016; Hsu et al., 2019b; Wang et al., 2018; Hsu et al., 2019a; Bensaid et al., 2021). Similarly, TaleBrush (Chung et al., 2022) leverages a canvas where sketching allows modulating the fortune of characters in a story, whereas Textoshop (Masson et al., 2025) offers features inspired by drawing software such as tone pickers and text layers.
Closest to our approach are systems that provide an editable alternative view of the text. For example, VISAR (Zhang et al., 2023) generates a visual outline of an argumentative text, which can be edited to change the structure of the argument. Toyteller (Chung et al., 2025) generates story text based on character motions, as if the characters were toys. Similarly, XCreation (Yan et al., 2023) supports creativity and cross-modality edits through the use of clip-arts and a relationship diagram as an intermediate representation between text and images.
@@
Our work contributes to this line of research by formalizing visual story-writing. The concept is rooted in the same principles as visual programming and code projections, positing that elements of a story become easier to understand and manipulate when presented in alternative forms. Specifically, we propose a framework grounded in structural narratology theories to inform the design of visualizations and extend the concept to all story elements, including events, like XCreation (Yan et al., 2023), but also space and time.
3. A Framework for Story Constructs
To inform the design of visual representations that support the visual story-writing workflow, we articulate a generative framework. Our framework builds upon narrative theories by structuralists like Gérard Genette (Genette et al., 1990) and Mieke Bal (Bal, 1997, 2021). Elements of stories outlined by structuralists (Bal, 1997, 2021) match what we observed in existing story visualizations, such as organizing story elements based on time (Munroe, 2009; Watson et al., 2019), space (Marti et al., 2018), or a combination thereof (Hulstein et al., 2023). Story elements do not work in isolation. Rather, many important aspects of stories stem from the intricate weaving of these elements, such as the movement of characters across locations over time.
We propose an operator mechanism to combine elements into such meaningful composites. Our operators are applied sequentially to story elements to form a story construct that can then be represented and manipulated. This framework can describe story constructs in existing visualizations and also generate new ones.
position to place elements spatially;
associate to add new elements;
connect to add edges; and
unfold to duplicate and organize elements.3.1. Story Elements
Story elements are the fundamental building blocks of story constructs found in story visualizations. Within narratology, the early Russian formalists separated the chronological order of the story (fabula) from the order of the plot or how it is represented to readers (syuzhet) (Liveley, 2021; Bal, 1997). Structuralists like Genette and Bal borrowed these concepts to analyze narrative structurally. Bal (Bal, 1997, 2021) describes a fabula, the factual events within a story, to consist of four main elements: actors111Actors refer to “agents that perform actions. They are not necessarily human” (Bal, 1997). To avoid confusion, we use the term entity instead in our prototype system., time, location, and events. The syuzhet, or how the story is told, also consists of four elements: characters, temporality, space, and focalization. See table 1 for examples.
| Element of Fabula | Element of Syuzhet |
| Actor (e.g., hero, villain) | Character (e.g., Alice) |
| Location (e.g., Alice’s house) | Space (e.g., home, eerie) |
| Time (chronological) | Temporality (narrated order) |
| Event (what happens) | Focalization (point of view) |
Although elements of fabula and syuzhet may seem similar, they refer to distinct concepts. For instance, characters are concrete entities (e.g., Alice, the white rabbit), whereas actors are abstract roles or functions characters can take (e.g., hero, villain). Time corresponds to the chronological timeline of events, whereas temporality corresponds to how those events are narrated (e.g., through flashbacks, ellipsis). And locations are concrete, physical places (e.g., Alice’s house, a forest), whereas spaces refer to the setting and the narrated representation of these locations (e.g., home, eerie). We describe visual story-writing as the use of visuals that represent and manipulate story constructs, i.e. composites of these eight elements.
3.2. Operators
Our operators are functions that combine story elements (operands) into higher-level story constructs. They can be chained to create more elaborate constructs. Based on existing story visualizations, we derived four binary operators acting on operands and 222Note that our operators are noncommutative. Thus, while is always one of the eight story elements, can be a composite of elements resulting from a chain of operations.:
position (
) places elements of based on the story element (Figure 2.
). For example, positioning time by locations results in a map of the movements in the story. Note that the only valid operands are location and space. -
associate (
) adds the elements of and associates them to the elements of (Figure 2.
For example, associating time with focalization creates a timeline of point of view. -
connect (
) adds edges between elements of , according to that of (Figure 2.
). For example, connecting characters by events creates a graph of character interactions. -
unfold (
) duplicates and organizes elements of based on that of (Figure 2.
). This essentially duplicates the elements according to each element of . For example, unfolding locations by characters creates a list of locations visited by each character.
3.3. Describing Existing Story Visualizations
We test our framework by describing existing story visualizations.
Storyline (Munroe, 2009) shows character interactions: start with time, unfold by characters, and connect by events.
StoryCurve (Kim et al., 2018) shows nonlinear narratives: start with time, unfold by temporality to visualize the non-linear narrative, associate with locations (using colours) and associate with events.
StoryPrint (Watson et al., 2019) shows scenes, character presence and emotions: start with time, unfold by characters, and associate with events or other elements (StoryPrint also encodes emotions, which goes beyond the scope of structural narratology).
Geo-Storylines (Hulstein et al., 2023) shows locations of characters over time: start with Storyline, position by locations. The map glyph variation: start with locations, position by location, unfold by time, and associate with characters.
“All fights from Dragon Ball Z”333A non-academic example (Cinnamon, 2017) shows the battles between characters within a saga: start with events (only the fights), unfold by time, and connect by characters.
3.4. Generating New Story Constructs
Figure 2 shows examples of chaining operators to create more complex story constructs. The number of unique combinations of operators and elements is vast. Our goal is not to argue which specific story construct is optimal for writers nor which visual design is best to represent these constructs—we leave it to future work to explore variations within these categories (and beyond). Instead, our work focuses on the potential of visual story-writing, i.e. the use of alternative representations of stories, as an input medium for expressing intent. As such, we designed three simple visual representations of story constructs interconnected through interactivity (North and Shneiderman, 2001; Wang Baldonado et al., 2000): a timeline of entities and events (start with time, unfold by characters and connect by events); a graph of actions and entities (start with characters and connect by events). And a view of the locations (start with characters and position by locations). There are other important components to a story, such as character traits and emotions, but in Bal’s narratology model (Bal, 1997, 2021), these are properties of the story elements. One way to represent them within the framework is as an attribute of a story element.
3.5. Recommending Interactions and Visuals
The framework can be further expanded to prescribe interactions and visual representations. For example, applying the position operator suggests that the placed elements can be dragged to express a change of location; the associate operator suggests a detail-on-demand type of interaction to view the associated elements; the connect operator suggests an interaction to remove and create links between elements; and the unfold operator suggests elements can be dragged to be reordered or reassigned. Similarly, the framework could also prescribe visual representations by mapping operations to specific visual channels.
4. The Visual Story-Writing Prototype
@@
To test the potential of visual story-writing, we developed a prototype system, shown in Figure 3, with three visual representations informed by our framework: a diagrammatic view of the entities and actions, a spatial view of the locations and entities, and a timeline of the events (section 3.4). The focus of the prototype was to help writers explore variations and perform spatial or temporal edits via visual story-writing. As such, the system allowed specifying edits through direct manipulations of the graphical elements (Shneiderman, 1982; Hutchins et al., 1985) and followed the design principles of creativity support tools to support exploration (Shneiderman et al., 2006). Below, we detail the system design and features (marked with
when the feature helps review the story,
when it helps edit the story, and
when it helps for both editing and reviewing). All screenshots were taken using the system on the first three paragraphs of the story Alice’s Adventures in Wonderland by Lewis Carroll.